### v1

# 3DGS.zip: A survey on 3D Gaussian Splatting Compression Methods

Abstract
We present a work-in-progress survey on 3D Gaussian Splatting[5] compression methods, focusing on their statistical performance across various benchmarks. This survey aims to facilitate comparability by summarizing key statistics of different compression approaches in a tabulated format. The datasets evaluated include TanksAndTemples[6], MipNeRF360[1], DeepBlending[4], and SyntheticNeRF[8]. For each method, we report the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Learned Perceptual Image Patch Similarity (LPIPS), and the resultant size in megabytes (MB), as provided by the respective authors. This is an ongoing, open project, and we invite contributions from the research community as GitHub issues or pull requests. Please visit http://w-m.github.io/3dgs-compression-survey/ for more information and a sortable version of the table.
Survey Table
Method TanksAndTemples MipNeRF360 DeepBlending SyntheticNeRF PSNR SSIM LPIPS Size MB PSNR SSIM LPIPS Size MB PSNR SSIM LPIPS Size MB PSNR SSIM LPIPS Size MB HAC-highrate 24.4 0.853 0.177 11.2 27.77 0.811 0.23 21.9 30.3 0.906 0.258 6.3 33.71 0.968 0.034 1.9 HAC-lowrate 24 0.846 0.187 8.1 27.5 0.807 0.238 15.3 30 0.902 0.269 4.3 33.2 0.967 0.037 1.2 LightGaussian 23.1 0.817 0.231 22 27.3 0.805 0.243 42 32.7 0.965 0.037 7.8 Compact3DGS 23.3 0.831 0.201 39.4 27.1 0.798 0.247 48.8 29.8 0.901 0.258 43.2 33.3 0.968 0.034 5.5 Compact3DGS+PP 23.3 0.831 0.202 20.9 27 0.797 0.247 29.1 29.7 0.9 0.258 23.8 Morgenstern et al. 25.63 0.864 0.208 21.4 27.6 0.814 0.22 40.3 30.4 0.909 0.258 16.8 33.7 0.969 0.031 4.1 Morgenstern et al. w/o SH 25.3 0.857 0.217 8.2 27 0.803 0.232 16.7 30.5 0.908 0.261 5.5 31.8 0.961 0.04 2 Navaneet et al. 16K 23.4 0.836 0.2 12 27 0.804 0.243 18 29.9 0.906 0.252 12 Navaneet et al. 32K 23.4 0.838 0.198 13 27.1 0.806 0.24 19 29.9 0.907 0.251 13 Compressed3D 23.3 0.832 0.194 17.3 27 0.801 0.238 28.8 29.4 0.898 0.253 25.3 32.9 0.967 0.033 3.7 Reduced3DGS 23.6 0.84 0.188 14 27.1 0.809 0.226 29 29.6 0.902 0.249 18
Contribution Summaries
The following sections provide brief summaries of the current contributions. For more detailed insights, please refer to the cited papers.
Reducing the Memory Footprint of 3D Gaussian Splatting
The authors reduce the memory requirements of 3D Gaussian Splatting for novel view synthesis by implementing an efficient resolution-aware primitive pruning approach, adaptive adjustment of spherical harmonics coefficients, and codebook-based quantization. The authors achieve a 27x reduction in memory size and a 1.7x increase in rendering speed with minimal impact on visual quality. [12]
LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS
LightGaussian aims to transform 3D Gaussians to a more efficient and compact form, avoiding the scalablity issues that arrises from the large number of SfM (Structure from Motion) points for unbounded scenes. Inspired by Network Pruning, the method identifies Gaussians that minimally contribute to scene reconstruction and employs a pruning and recovery process, thereby efficiently reducing redundancy in Gaussian counts while maintaining visual effects. Additionally, LightGaussian utilizes knowledge distillation and pseudo-view augmentation to transfer spherical harmonics efficients to a lower degree. Furthermore, the authors propose a Gaussian Vector Quantization based on the global significance of Gaussians to quantize all redundant attributes, achieving lower bitwidth representations with minimal accuracy losses. [3]
HAC: Hash-grid Assisted Context for 3D Gaussian Splatting Compression
The paper proposes a Hash-grid Assisted Context (HAC) framework for compressing 3D Gaussian Splatting (3DGS) models by leveraging the mutual information between attributes of unorganized 3D Gaussians (anchors) and hash grid features. Using Scaffold-GS as a base model, HAC queries the hash grid by anchor location to predict anchor attribute distributions for efficient entropy coding. The framework introduces an Adaptive Quantization Module (AQM) to dynamically adjust quantization step sizes and employs adaptive offset masking with learnable masks to prune unnecessary data for further compression. Extensive experiments show that HAC achieves significant compression ratios while maintaining or improving rendering fidelity. [2]
Compact 3D Gaussian Representation for Radiance Field
This approach aims to reduce the number of Gaussian points neccecary for 3D Gaussian splat representations without sacrificing performance, and to compresse the Gaussian attributes, such as view-dependent color and covariance. The authors observe current densification algorithm to produces many redundant and unnecessary Gaussians, resulting in high memory and storage requirements. Thus, they introduce a novel volume-based masking strategy that identifies and removes non-essential Gaussians that have minimal impact on overall performance. Furthermore, they exploit spatial redundancy by incorporating a grid-based (Instant-NGP) neural field to efficiently represent view-dependent colors rather than using per Gaussian color attributes. Given that the geometry of the Gaussians is expected to have limited variation in scale and rotation, this approach employs a compact codebook-based representation, which is designed to identify similar patterns or geometries shared across each scene. [7]
Compact3D: Compressing Gaussian Splat Radiance Field Models with Vector Quantization
This approach introduces a vector quantization method based on the K-means algorithm to quantize the Gaussian parameters in 3D Gaussian splatting, as many Gaussians may share similar parmeters. Only a small codebook is stored along with the index of the code for each Gaussian, resulting in a large reduction in the storage of the learned radiance fields and a reduction of the memory footprint at rendering time. Additionally, the representation is further compressed by sorting the Gaussians based on one of the quantized parameters and storing the indices using the Run-Length-Encoding (RLE) method. [10]
Compact 3D Scene Representation via Self-Organizing Gaussian Grids
Compressing 3D data is challenging, but many effective solutions exist for compressing 2D data (such as images). The authors propose a new method to organize 3DGS parameters into a 2D grid, drastically reducing storage requirements without compromising visual quality. This organization exploits perceptual redundancies in natural scenes. They introduce a highly parallel sorting algorithm, PLAS, which arranges Gaussian parameters into a 2D grid, maintaining local neighborhood structure and ensuring smoothness. This solution is particularly innovative because no existing method efficiently handles a 2D grid with millions of points. During training, a smoothness loss is applied to enforce local smoothness in the 2D grid, enhancing the compressibility of the data. The key insight is that smoothness needs to be enforced during training to enable efficient compression. [9]
Compressed 3D Gaussian Splatting for Accelerated Novel View Synthesis
This approach aims to reduce the memory consumption and rendering efficency of 3D Gaussian splat representations. The authors propose a compressed 3D Gaussian splat representation consisting of three main steps: 1. sensitivity-aware clustering, where scene parameters are measured according to their contribution to the training images and encoded into compact codebooks via sensitivity-aware vector quantization; 2. quantization-aware fine-tuning, which recovers lost information by fine-tuning parameters at reduced bit-rates using quantization-aware training; and 3. entropy encoding, which exploits spatial coherence through entropy and run-length encoding by linearizing 3D Gaussians along a space-filling curve. Furthermore, a renderer for the compressed scenes utilizing GPU-based sorting and rasterization is proposed, enabling real-time novel view synthesis on low-end devices. [11]