### v2

# 3DGS.zip: A survey on 3D Gaussian Splatting Compression Methods

Abstract
We present a work-in-progress survey on 3D Gaussian Splatting[5] compression methods, focusing on their statistical performance across various benchmarks. This survey aims to facilitate comparability by summarizing key statistics of different compression approaches in a tabulated format. The datasets evaluated include TanksAndTemples[6], MipNeRF360[1], DeepBlending[4], and SyntheticNeRF[8]. For each method, we report the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Learned Perceptual Image Patch Similarity (LPIPS), and the resultant size in megabytes (MB), as provided by the respective authors. This is an ongoing, open project, and we invite contributions from the research community as GitHub issues or pull requests. Please visit http://w-m.github.io/3dgs-compression-survey/ for more information and a sortable version of the table.
Survey Table
Method
Rank
TanksAndTemples
MipNeRF360
DeepBlending
SyntheticNeRF
PSNR
SSIM
LPIPS
Size
MB
PSNR
SSIM
LPIPS
Size
MB
PSNR
SSIM
LPIPS
Size
MB
PSNR
SSIM
LPIPS
Size
MB
HAC-lowrate
3.0
24.04
0.846
0.187
8.1
27.53
0.807
0.238
15.3
29.98
0.902
0.269
4.3
33.24
0.967
0.037
1.2
HAC-highrate
3.2
24.40
0.853
0.177
11.2
27.77
0.811
0.230
21.9
30.34
0.906
0.258
6.3
33.71
0.968
0.034
1.9
LightGaussian
8.9
23.11
0.817
0.231
22.0
27.28
0.805
0.243
42.0
32.725
0.965
0.037
7.8
EAGLES
8.8
23.37
0.84
0.20
29.0
27.23
0.81
0.24
54.0
29.86
0.91
0.25
52.0
EAGLES-Small
9.9
23.1
0.82
0.22
19.0
26.94
0.80
0.25
47.0
29.92
0.90
0.25
33.0
Compact3DGS
9.2
23.32
0.831
0.201
39.4
27.08
0.798
0.247
48.8
29.79
0.901
0.258
43.2
33.33
0.968
0.034
5.5
Compact3DGS+PP
9.3
23.32
0.831
0.202
20.9
27.03
0.797
0.247
29.1
29.73
0.900
0.258
23.8
Morgenstern et al.
5.1
25.63
0.864
0.208
21.4
27.64
0.814
0.220
40.3
30.35
0.909
0.258
16.8
33.70
0.969
0.031
4.1
Morgenstern et al. w/o SH
4.2
25.27
0.857
0.217
8.2
27.02
0.803
0.232
16.7
30.50
0.908
0.261
5.5
31.75
0.961
0.040
2.0
Navaneet et al. 16K
5.5
23.39
0.836
0.200
12.0
27.03
0.804
0.243
18.0
29.90
0.906
0.252
12.0
Navaneet et al. 32K
5.2
23.44
0.838
0.198
13.0
27.12
0.806
0.240
19.0
29.90
0.907
0.251
13.0
Compressed3D
7.0
23.324
0.832
0.194
17.3
26.981
0.801
0.238
28.8
29.381
0.898
0.253
25.3
32.936
0.967
0.033
3.7
Reduced3DGS
5.9
23.57
0.84
0.188
14.0
27.1
0.809
0.226
29.0
29.63
0.902
0.249
18.0
Note: The best methods in each category is highlighted. The ranks represent the average rankings of the methods across all available datasets. The quality metrics PSNR, SSIM, and LPIPS are equally weighted with the model size, meaning they each contribute one-sixth to the ranks, while the size contributes half.
Contribution Summaries
The following sections provide brief summaries of the current contributions. For more detailed insights, please refer to the cited papers.
Reducing the Memory Footprint of 3D Gaussian Splatting
The authors reduce the memory requirements of 3D Gaussian Splatting for novel view synthesis by implementing an efficient resolution-aware primitive pruning approach, adaptive adjustment of spherical harmonics coefficients, and codebook-based quantization. The authors achieve a 27x reduction in memory size and a 1.7x increase in rendering speed with minimal impact on visual quality. [12]
LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS
LightGaussian aims to transform 3D Gaussians to a more efficient and compact form, avoiding the scalablity issues that arrises from the large number of SfM (Structure from Motion) points for unbounded scenes. Inspired by Network Pruning, the method identifies Gaussians that minimally contribute to scene reconstruction and employs a pruning and recovery process, thereby efficiently reducing redundancy in Gaussian counts while maintaining visual effects. Additionally, LightGaussian utilizes knowledge distillation and pseudo-view augmentation to transfer spherical harmonics efficients to a lower degree. Furthermore, the authors propose a Gaussian Vector Quantization based on the global significance of Gaussians to quantize all redundant attributes, achieving lower bitwidth representations with minimal accuracy losses. [3]
HAC: Hash-grid Assisted Context for 3D Gaussian Splatting Compression
The paper proposes a Hash-grid Assisted Context (HAC) framework for compressing 3D Gaussian Splatting (3DGS) models by leveraging the mutual information between attributes of unorganized 3D Gaussians (anchors) and hash grid features. Using Scaffold-GS as a base model, HAC queries the hash grid by anchor location to predict anchor attribute distributions for efficient entropy coding. The framework introduces an Adaptive Quantization Module (AQM) to dynamically adjust quantization step sizes and employs adaptive offset masking with learnable masks to prune unnecessary data for further compression. Extensive experiments show that HAC achieves significant compression ratios while maintaining or improving rendering fidelity. [2]
EAGLES: Efficient Accelerated 3D Gaussians with Lightweight EncodingS
This approach aims to decrease both storage and runtime memory costs while enhancing training and rendering speeds and maintaining view synthesis quality. The authors observed that the color and rotation attributes account for over 80
Compact 3D Gaussian Representation for Radiance Field
This approach aims to reduce the number of Gaussian points neccecary for 3D Gaussian splat representations without sacrificing performance, and to compresse the Gaussian attributes, such as view-dependent color and covariance. The authors observe current densification algorithm to produces many redundant and unnecessary Gaussians, resulting in high memory and storage requirements. Thus, they introduce a novel volume-based masking strategy that identifies and removes non-essential Gaussians that have minimal impact on overall performance. Furthermore, they exploit spatial redundancy by incorporating a grid-based (Instant-NGP) neural field to efficiently represent view-dependent colors rather than using per Gaussian color attributes. Given that the geometry of the Gaussians is expected to have limited variation in scale and rotation, this approach employs a compact codebook-based representation, which is designed to identify similar patterns or geometries shared across each scene. [7]
Compact3D: Compressing Gaussian Splat Radiance Field Models with Vector Quantization
This approach introduces a vector quantization method based on the K-means algorithm to quantize the Gaussian parameters in 3D Gaussian splatting, as many Gaussians may share similar parmeters. Only a small codebook is stored along with the index of the code for each Gaussian, resulting in a large reduction in the storage of the learned radiance fields and a reduction of the memory footprint at rendering time. Additionally, the representation is further compressed by sorting the Gaussians based on one of the quantized parameters and storing the indices using the Run-Length-Encoding (RLE) method. [10]
Compact 3D Scene Representation via Self-Organizing Gaussian Grids
Compressing 3D data is challenging, but many effective solutions exist for compressing 2D data (such as images). The authors propose a new method to organize 3DGS parameters into a 2D grid, drastically reducing storage requirements without compromising visual quality. This organization exploits perceptual redundancies in natural scenes. They introduce a highly parallel sorting algorithm, PLAS, which arranges Gaussian parameters into a 2D grid, maintaining local neighborhood structure and ensuring smoothness. This solution is particularly innovative because no existing method efficiently handles a 2D grid with millions of points. During training, a smoothness loss is applied to enforce local smoothness in the 2D grid, enhancing the compressibility of the data. The key insight is that smoothness needs to be enforced during training to enable efficient compression. [9]
Compressed 3D Gaussian Splatting for Accelerated Novel View Synthesis
This approach aims to reduce the memory consumption and rendering efficency of 3D Gaussian splat representations. The authors propose a compressed 3D Gaussian splat representation consisting of three main steps: 1. sensitivity-aware clustering, where scene parameters are measured according to their contribution to the training images and encoded into compact codebooks via sensitivity-aware vector quantization; 2. quantization-aware fine-tuning, which recovers lost information by fine-tuning parameters at reduced bit-rates using quantization-aware training; and 3. entropy encoding, which exploits spatial coherence through entropy and run-length encoding by linearizing 3D Gaussians along a space-filling curve. Furthermore, a renderer for the compressed scenes utilizing GPU-based sorting and rasterization is proposed, enabling real-time novel view synthesis on low-end devices. [11]
F-3DGS: Factorized Coordinates and Representations for 3D Gaussian Splatting
The paper introduces a novel 3D Gaussian compression method using structured coordinates and decomposed representations through factorization. Inspired by tensor or matrix factorization techniques, this method generates 3D coordinates via a tensor product of 1D or 2D coordinates, enhancing spatial efficiency. It extends factorization to include attributes like color, scale, rotation, and opacity, which compresses the model size while maintaining essential characteristics. A binary mask is employed to eliminate non-essential Gaussians, significantly accelerating training and rendering speeds. Experiments demonstrate that this approach reduces storage requirements by over 90<br><br> <span style="color:red"><b>Note:</b> This paper is currently not included in the survey table because it shows unusually high results in the Tanks and Temples dataset and reports higher results for the original 3DGS than those in the original publication. This raises the possibility that their testing methods may differ from those used in other papers.</span> [13]